# Exploring Vacant Classes in Label-Skewed Federated Learning

### v2

@@ Abstract
Label skews, characterized by disparities in local label distribution across clients, pose a significant challenge in federated learning. As minority classes suffer from worse accuracy due to overfitting on local imbalanced data, prior methods often incorporate class-balanced learning techniques during local training. Although these methods improve the mean accuracy across all classes, we observe that vacant classes—referring to categories absent from a client’s data distribution—remain poorly recognized. Besides, there is still a gap in the accuracy of local models on minority classes compared to the global model. This paper introduces FedVLS, a novel approach to label-skewed federated learning that integrates both vacant-class distillation and logit suppression simultaneously. Specifically, vacant-class distillation leverages knowledge distillation during local training on each client to retain essential information related to vacant classes from the global model. Moreover, logit suppression directly penalizes network logits for non-label classes, effectively addressing misclassifications in minority classes that may be biased toward majority classes. Extensive experiments validate the efficacy of FedVLS, demonstrating superior performance compared to previous state-of-the-art (SOTA) methods across diverse datasets with varying degrees of label skews. Code is available in the supplementary material.
@@ Introduction
Figure 1: Class-wise accuracy of the initial global model and updated local model on IID and label-skewed CIFAR10 data distributions. (a) represents the result updating on IID local data with FedAvg (McMahan et al. 2017). (b-d) showcase the results updating on skewed local data distribution with FedAvg (McMahan et al. 2017), FedLC (Zhang et al. 2022), and our FedVLS, respectively. The value (%) in each caption corresponds to the accuracy of the global model aggregated from updated local models.
@@
  • We find that prior federated learning methods suffer from the existence of vacant classes and propose FedVLS to distill vacant-class-aware knowledge from the global model.
@@
Federated learning faces a significant challenge known as data heterogeneity, also referred to as non-identical and independently distributed (Non-IID) data (Kairouz et al. 2021; Luo et al. 2021; Shi et al. 2023b). This challenge encompasses issues such as label skews and domain shifts. In this paper, we primarily focus on addressing label skews. The classic federated learning algorithm, FedAvg (McMahan et al. 2017), experiences a significant decline in performance when dealing with label skews (Li et al. 2019; Acar et al. 2021). Numerous studies have aimed to mitigate the adverse impacts of label skews. For instance, FedProx (Li et al. 2020b) employs a proximal term and SCAFFOLD (Karimireddy et al. 2020) uses a variance reduction approach to constrain the update direction of local models. Additionally, MOON (Li, He, and Song 2021) and FedProc (Mu et al. 2023) utilize contrastive loss to enhance the agreement between local models and the global model. Furthermore, FedConcat (Diao, Li, and He 2024) propose model concatenation, FedMR (Fan et al. 2023) proposes a manifold reshaping approach, FedGELA (Fan et al. 2024) uses simplex Equiangular Tight to initialize the local classifier and FedGF (Lee and Yoon 2024) refine the flat minima searching to alleviate the label skews. However, these methods often fail to address the issue of vacant classes in highly skewed scenarios. We propose FedVLS to effectively mitigate the decline in class-wise accuracy of vacant classes.
@@
Unlike FedNTD (Lee et al. 2022), which encourages the client model to closely match the global model’s output for not-true labels, thereby limiting the knowledge protection for vacant class, our loss function ensures that the local model replicates the global model’s outputs only for vacant-class labels. This approach preserves the predictive capability for vacant categories significantly. Moreover, the computational overhead introduced by this loss function is minimal, enhancing its practical implementation. Additional comparisons with other distillation-based methods and further analyses are provided in the technical appendix
@@
Implementation details We set the number of clients N to 10 and implement full client participation. We run 100 communication rounds for all experiments on the CIFAR10/100 datasets and 50 communication rounds on the MNIST and TinyImageNet datasets. Within each communication round, local training spans 5 epochs for MINST and 10 epochs for the other datasets. For FedConcat (Diao, Li, and He 2024) and FedGF (Lee and Yoon 2024), we followed the original paper’s settings for communication rounds and local epochs. We employ stochastic gradient descent (SGD) optimization with a learning rate of 0.01, a momentum of 0.9, and a batch size of 64. Weight decay is set to $10^{-5}$ for MNIST and CIFAR10 and $10^{-4}$ for CIFAR100 and TinyImageNet. The hyperparameter $\lambda$ of FedVLS in Equation 6 is set to 0.1 for MNIST and CIFAR10, while it is set to 0.5 for CIFAR100 and TinyImageNet. Following pFedMe (T Dinh, Tran, and Nguyen 2020), we conduct three trials for each experimental setting and report the mean accuracy and standard deviation of the maximum accuracy achieved by the global model during the training process. More implementation details can be found in the technical appendix.

### v3

@@ Abstract
Label skews, characterized by disparities in local label distribution across clients, pose a significant challenge in federated learning. As minority classes suffer from worse accuracy due to overfitting on local imbalanced data, prior methods often incorporate class-balanced learning techniques during local training. Although these methods improve the mean accuracy across all classes, we observe that vacant classes—referring to categories absent from a client’s data distribution—remain poorly recognized. Besides, there is still a gap in the accuracy of local models on minority classes compared to the global model. This paper introduces FedVLS, a novel approach to label-skewed federated learning that integrates both vacant-class distillation and logit suppression simultaneously. Specifically, vacant-class distillation leverages knowledge distillation during local training on each client to retain essential information related to vacant classes from the global model. Moreover, logit suppression directly penalizes network logits for non-label classes, effectively addressing misclassifications in minority classes that may be biased toward majority classes. Extensive experiments validate the efficacy of FedVLS, demonstrating superior performance compared to previous state-of-the-art (SOTA) methods across diverse datasets with varying degrees of label skews. Our code is available at https://github.com/krumpguo/FedVLS.
@@ Introduction
Figure 1: Class-wise accuracy of the initial global model and updated local modelS on IID and label-skewed CIFAR10 data distributions. (a) represents the result updating on IID local data with FedAvg (McMahan et al. 2017). (b-d) showcase the results updating on skewed data distribution with FedAvg (McMahan et al. 2017), FedLC (Zhang et al. 2022), and our FedVLS, respectively. The value (%) in each caption corresponds to the accuracy of the global model aggregated from local models.
@@
  • We find that prior federated learning methods suffer from vacant classes and propose FedVLS to distill vacant-class-aware knowledge from the global model.
@@
Federated learning faces a significant challenge known as data heterogeneity, also referred to as non-identical and independently distributed (Non-IID) data (Kairouz et al. 2021; Luo et al. 2021; Shi et al. 2023b; Guo, Wang, and Geng 2024; Guo et al. 2024). This challenge encompasses issues such as label skews and domain shifts. In this paper, we primarily focus on addressing label skews. The classic federated learning algorithm, FedAvg (McMahan et al. 2017), experiences a significant decline in performance when dealing with label skews (Li et al. 2019; Acar et al. 2021; Luo, Wang, and Wang 2024). Numerous studies have aimed to mitigate the adverse impacts of label skews. For instance, FedProx (Li et al. 2020b) employs a proximal term and SCAFFOLD (Karimireddy et al. 2020) uses a variance reduction approach to constrain the update direction of local models. Additionally, MOON (Li, He, and Song 2021) and FedProc (Mu et al. 2023) utilize contrastive loss to enhance the agreement between local models and the global model. Furthermore, FedConcat (Diao, Li, and He 2024) propose model concatenation, FedMR (Fan et al. 2023) proposes a manifold reshaping approach, FedGELA (Fan et al. 2024) uses simplex Equiangular Tight to initialize the local classifier and FedGF (Lee and Yoon 2024) refine the flat minima searching to alleviate the label skews. However, these methods often fail to address the issue of vacant classes in highly skewed scenarios. We propose FedVLS to effectively mitigate the decline in class-wise accuracy of vacant classes.
@@
Unlike FedNTD (Lee et al. 2022), which encourages the client model to closely match the global model’s output for not-true labels, thereby limiting the knowledge protection for vacant class, our loss function ensures that the local model replicates the global model’s outputs only for vacant-class labels. This approach preserves the predictive capability for vacant categories significantly. Moreover, the computational overhead introduced by this loss function is minimal, enhancing its practical implementation. Additional comparisons with other distillation-based methods and further analyses are provided in the technical appendix.
@@
Implementation details We set the number of clients N to 10 and implement full client participation. We run 100 communication rounds for all experiments on the CIFAR10/100 datasets and 50 communication rounds on the MNIST and TinyImageNet datasets. Within each communication round, local training spans 5 epochs for MNIST and 10 epochs for the other datasets. For FedConcat (Diao, Li, and He 2024) and FedGF (Lee and Yoon 2024), we followed the original paper’s settings for communication rounds and local epochs. We employ stochastic gradient descent (SGD) optimization with a learning rate of 0.01, a momentum of 0.9, and a batch size of 64. Weight decay is set to $10^{-5}$ for MNIST and CIFAR10 and $10^{-4}$ for CIFAR100 and TinyImageNet. The hyperparameter $\lambda$ of FedVLS in Equation 6 is set to 0.1 for MNIST and CIFAR10, while it is set to 0.5 for CIFAR100 and TinyImageNet. Following pFedMe (T Dinh, Tran, and Nguyen 2020), we conduct three trials for each experimental setting and report the mean accuracy and standard deviation of the maximum accuracy achieved by the global model during the training process. More implementation details and experimental results can be found in the technical appendix at https://github.com/krumpguo/FedVLS.
@@ Conclusion
Acknowledgments
This work was funded by the National Natural Science Foundation of China under Grants (62276256, U2441251) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).